OpenStudy (anonymous):

Earl opens a savings account with $5,000. He deposits $1500 each year into the account that compounds quarterly and has a 0.65% interest rate. What will his account total be in 5 years? $14,331.46 $12,781.95 $35,632.71 $15,681.02

OpenStudy (dumbcow):

its a geometric series \[B = 5,000(1+i)^{20}+1500[1+(1+i)^{4}+(1+i)^{8}+(1+i)^{12}+(1+i)^{16}]\] assuming he make deposit at end of year i = .0065/4 use formula for sum of geometric series \[B = 5,000(1+i)^{20} +1500\frac{(1+i)^{20}-1}{(1+i)-1}\]

OpenStudy (anonymous):

i bet everyones head is about to explode right now

OpenStudy (anonymous):

what i

OpenStudy (anonymous):

what's i

OpenStudy (anonymous):

whoops wasnt paying attention ok

OpenStudy (anonymous):

i didnt get the right answer

OpenStudy (dumbcow):

oh theres an error somewhere...the bottom of fraction should be (1+i)^4 -1 since thats whats being multiplied each time

OpenStudy (dumbcow):

still don't get their answer exactly...buts its pretty close to B

OpenStudy (anonymous):

yeah mine came close to b too.

OpenStudy (dumbcow):

also since this is multiple choice, its easy....imagine you earned no interest --> 5000 + 1500*5 = 12,500 0.65% interest is really small so the earnings will be small , this leaves only 1 possible answer

OpenStudy (anonymous):

okay, thanks for helping dumbcow :))